OpenStudy (anonymous):

Find the diameter of a sphere with surface area 401.7cm2.

OpenStudy (saifoo.khan):

Do you know the formula for the surface area of a sphere?

OpenStudy (saifoo.khan):

http://0.tqn.com/d/math/1/0/I/F/spherer.gif

OpenStudy (anonymous):

SA=4pi*r^2

OpenStudy (anonymous):

i plugged in the SA=401.7/4pi, the r^2=315.499....i am unsure if i approach the question correctly

OpenStudy (saifoo.khan):

correct.

OpenStudy (anonymous):

you are right @1dpp

OpenStudy (anonymous):

the diameter is doubled, but as i entered my answer, i incorrect?

OpenStudy (anonymous):

\[401.7 = {4\times3.14\times r^2}\] now solve for r

OpenStudy (anonymous):

the answer =11 diameter....thanks :/

OpenStudy (anonymous):

it should be 11.31 cm